If three mangoes are taken one after the other from a basket which contain 10 ripe and 4 unripe


mangoes



1. What is the probability that you will get 3 ripe mangoes?


2. What is the probability that you will get 2 ripe mangoes?


3. What is the probability that you will get 1 ripe mango?


4. What is the probability that you will not get ripe mango?

1
Expert's answer
2022-03-01T18:52:29-0500

We can take three mangoes in "\dbinom{10+4}{3}=364" possible ways.

1.

"P(3\ ripe)=\dfrac{\dbinom{10}{3}\dbinom{4}{3-3}}{\dbinom{14}{3}}=\dfrac{120(1)}{364}=\dfrac{30}{91}"

2.

"P(2\ ripe\ \& \ 1\ unripe)=\dfrac{\dbinom{10}{2}\dbinom{4}{3-2}}{\dbinom{14}{3}}=\dfrac{45(4)}{364}=\dfrac{45}{91}"

3.

"P(1\ ripe\ \& \ 2\ unripe)=\dfrac{\dbinom{10}{1}\dbinom{4}{3-1}}{\dbinom{14}{3}}=\dfrac{10(6)}{364}=\dfrac{15}{91}"



4.

"P(3\ unripe)=\dfrac{\dbinom{10}{0}\dbinom{4}{3-0}}{\dbinom{14}{3}}=\dfrac{1(4)}{364}=\dfrac{1}{91}"